The magnetic eld of an electric charge q is given by B =

0

4

qv sin 

r2 . From this equation, we can

deduce that

(a) The magnetic eld is only created by a moving charge.

(b) If the angle  is zero, then the magnetic eld is also zero. [3]

(c) The magnetic eld decrease as r decreases.

(d) If the charge increases, the magnetic eld decreases.

(e) None of the above.


1
Expert's answer
2021-11-04T18:05:19-0400

Now

"B=\frac{\mu_0}{4\pi}\frac{qvsin\theta}{r^2}"

Option (c) is correct option

The magnetic field decrease by the r is decreased is incorrect statement

But magnetic field inverse perportional to squre of distance

"B\alpha \frac{1}{r^2}"


"r increase ; B decrease"

Option (c) is false statements